Your Composition 2 professor is consistently between 2 and 13 minutes late for class, following a uniform probability distribution. What is the probability that they will be between 4 and 11 minutes late today? Round to 4 decimal places, as needed.
Answer #1

Solution :

Given that,

a = 2

b = 13

P(c < x < d) = (d - c) / (b - a)

P(4 < x < 11) = (11 - 4) / (13 - 2)

= 7 / 11 = 0.6364

Probability = 0.6364